a geography test has 6 questions worth 3 marks each and 16 questions worth 2 marks each. Write an algebraic expression for each situation below.
1
Expert's answer
2020-03-31T15:59:05-0400

1. The total score for the test 6 * 3 + 16 * 2 = 18 + 32 = 50

2. The total score for questions of 3 points 6 * 3 = 18

3. The total score for questions of 2 points 16 * 2 = 32